OpenStudy (anonymous):

Solve the following system of equations and show all work. y = −x^2 + 4 y = 2x + 1

OpenStudy (anonymous):

i know how to solve a normal system of equations but the squared x is confusing me

OpenStudy (haseeb96):

consider equation 1 as eq#1 and equation 2 a eq#2 y= -x^2+4 .............eq#1 y=2x+1...................eq#2 now just pput this eq#2 into eq#1 at the place of y

OpenStudy (anonymous):

2x+1=x^2+4

OpenStudy (haseeb96):

no 2x+1 = -x^2+4 x^2+2x-3 =0 now factorize this equation and tell me what you get the value of x ?

OpenStudy (haseeb96):

do you know how to do factorization?

OpenStudy (anonymous):

oohh yes of course you'd get x=-3,1

OpenStudy (anonymous):

thx lots

OpenStudy (haseeb96):

sorry i was wrong @gold.shosho you were correct now plug this x values one by one in eq#2 and get the values of y

OpenStudy (nincompoop):

so how is the x^2 confusing?

OpenStudy (nincompoop):